OpenStudy (19rutlehop):

The video games store just increased the price of renting a game by $1.25 Question - Write a formula that will help Seth determine how much he will spend at the video store after the prices increase? p=old price of a video game n= Number of games Seth rents t= Total amount Seth pays

OpenStudy (priyar):

what will be the new price of 1 vido game?

OpenStudy (19rutlehop):

It's not telling me that, it's about Literal Equations

OpenStudy (sagewilson):

So normally without the added price it would look like this P * N = T but now that you added 1.25 to the P how would the equation look?

OpenStudy (19rutlehop):

Well thanks you P*N=T was the answer.

OpenStudy (sagewilson):

oh I thought it would have been (P +1.25) * N = T

OpenStudy (sagewilson):

But hey you got the answer :)

OpenStudy (priyar):

hey thats not the answer...what @SageWilson wrote later is correct